OpenStudy (mitchi35):

3) Graph the circle x2+ (y + 2)2= 9. Label the center and at least four points on the circle.

OpenStudy (mitchi35):

Number 3 please? Thanks!

OpenStudy (loser66):

what is the standard form of a circle?

OpenStudy (mitchi35):

(x-h)^2+(y-v)^2=r2

OpenStudy (loser66):

now compare to your expression, \(x^2 +(y+2)^2=9\) , then h =? k =? r =?

OpenStudy (mitchi35):

Okay....Wait. But How can I figure out the center points?

OpenStudy (loser66):

Why don't you answer me?:

OpenStudy (mitchi35):

Well geez. Alright.

OpenStudy (mitchi35):

I don't even understand what you are asking me to do.

OpenStudy (loser66):

Since you skip it, you don't know how to find the center. I realize that fact right after you ask me what is the center although you said ok before.

OpenStudy (mitchi35):

Not with this question no.

OpenStudy (loser66):

ok, slower. Do you agree with me that \(x^2 = (x-0)^2\)?

OpenStudy (mitchi35):

Yeah I guess

OpenStudy (mitchi35):

Yes I do.

OpenStudy (loser66):

and \((y +2) = y-(-2) \) , ok?

OpenStudy (mitchi35):

Yes

OpenStudy (loser66):

9 = 3^2 ok?

OpenStudy (mitchi35):

Ok

OpenStudy (loser66):

Compare the standard form of a circle with your problem: \[(x -\color{red}{h})^2 +(y -\color{blue}{k})^2 =\color{purple}{r}^2\\(x -\color{red}{0})^2 +(y -\color{blue}{(-2)})^2 =\color{purple}{3}^2\] Can you tell me h, k,r on your problem are?

OpenStudy (mitchi35):

Oh! Omg! Nevermind! I have it figured out. It's right here in my book. I had to dig to find it. But thank you! Do you want me to share it with you anyway?